# Modelling vehicle and pedestrian collective dynamics: Challenges and advancements

### v1

@@ 2.2.3. Hybrid models
Many pedestrian models are hybrid approaches combining different concepts and mechanisms from first and second order models such as social force, optimal velocity (OV), or optimisation and collision avoidance techniques. For example, some extensions of force-based models are inspired by swarming approaches [48, 66], where the desired velocity $\vec{v}_{0}$ in (11) is mitigated with the average velocity of the neighbouring pedestrians using
@@
to address mass collective behaviour. In the same vein, the pedestrian velocity is relaxed to a combination of the desired velocity and a superposition of distance-based OV models [87, 88], generalising concepts of car-following models [6, 5, 57] to two dimensions
@@
Another hybrid two-dimensional OV pedestrian model relies on first-order dynamics [120]. As in the modelling framework introduced by Maury and Venel in [80, 81], this model constrains pedestrian velocities to remain in admissible collision-free configurations [120]. However, the speed is regulated according to the distance to the next pedestrian in the direction of motion using an OV function, while the direction results from a superposition of distance-based repulsion with the neighbours as in force-based models. An extension of this collision-free velocity-based model relaxes the direction, making it a second-order model [143]. In the Gradient Navigation model introduced in [31], the speed is relaxed to the OV function, while the direction results from a the desired velocity and a superposition of distance-based repulsion with the neighbours as in force-based models
@@
where $C_{1}$ and $C_{2}$ are two normalisation constants while $V(\rho_{n})$ is the OV depending on the local density $\rho_{n}$. A similar approach is given in [86] with an heuristic model without social interaction and where, as in the collision-free velocity-based model [120], the OV depends on the distance to the nearest pedestrian in front.
The concept of anticipatory collision avoidance in force-based models introduced in [145] is deepened in a recent hybrid force-based model [58], where the interaction force is derived directly from the potential
@@
The TTC $\tau_{n,m}$ is the time before a collision occurs between the $n$-th and $m$-th pedestrians (i.e., $|\Delta\dot{\vec{x}}_{n,m}|<\ell$), assuming that their velocities are constant. In [74], the force is based on similar collision anticipation mechanisms. A general hybrid modelling framework first introduced in the early 2000s [53] assumes pedestrian dynamics based on a decisional layer, resulting from an optimal control problem, and a mechanical layer modelling physical exclusion force and other inertia phenomena
(25) $\tilde{\vec{v}}_{0}^{n}(t)=\text{arg}\min_{\vec{v}_{n}\in\mathbb{R}^{2}}\mathcal{E}_{\vec{x}(t)}^{\vec{v}_{0}}(\vec{v}_{n}),\qquad\ddot{\vec{x}}_{n}(t)=\frac{1}{\tau}\big{[}\tilde{\vec{v}}_{0}^{n}(t)-\dot{\vec{x}}_{n}(t)\big{]}+F^{\text{M}}_{n}(t),$
where $\mathcal{E}$ is an energy function involving several concepts, while $F^{\text{M}}_{n}(t)$ are the mechanical force on the $n$-th pedestrian at time $t$. In the ANticipatory Dynamic Algorithms (ANDA) [34], the energy function consists of four main components:
@@
    Anticipation energy based on the TTC potential (23)-(24) introduced in [58].
@@ 3. Modelling collective dynamics
Traffic and pedestrian flows describe many coordinated collective dynamics [46, 19, 16, 10] A typical example in traffic flow is stop-and-go dynamics on highways where jam waves resulting from instability phenomena [96, 139], also called jamitons [38], propagate upstream in the flow without any apparent external cause. Typical examples for pedestrian dynamics are lane formation for counterflow, jamming and clogging at bottlenecks, oscillation and flow intermittency for counterflow at bottlenecks, or stripe formation for crossing flow [47, 10]. Most of the aforementioned models describe these phenomena, with realistic features for fine-tuning the parameters. However, conventional models fall short of replicating more complex situations involving long-term anticipation (on the order of several seconds) or socio-psychological aspects. A typical example is the formation of a clear lane in pedestrian crowds to facilitate the passage of a vehicle, typically an ambulance [2], which relies on early anticipation. Another example in the load balancing of pedestrians over several exits in case of normal evacuation. Here, the distance between the exits requires pedestrians to anticipate for long periods of time. These different collective phenomena and related modelling approaches are discussed in more detail in the following sections.
@@ 3.1.1. Empirics of stop-and-go waves in traffic flow
Stop-and-go dynamics are currently observed in road traffic flow. They are also found in single-file pedestrian movements and in bicycle flow [147]. In fact, stop-and-go and accordion traffic now commonly denotes congested traffic. In addition to its scientific interest, the jerky dynamics of congested traffic is a major safety concern. It also has a negative impact on the environment. Indeed, the acceleration and deceleration phases of stop-and-go traffic dynamics cause an increase in fuel consumption and pollutant emissions [71].
The terminology of stop-and-go waves can be traced back to the late 1960s [32]. A related word is phantom jam [114], as the waves seem to appear for no apparent reason. This was demonstrated in the late 2000s with an experiment involving 22 vehicles on a 230-metre circuit [114] (see Fig. 3). Starting from a homogeneous configuration with equal distance gaps, a stop-and-go wave emerged after a while, with jerky vehicular dynamics not caused by any infrastructure peculiarity. This experiment shed new light on the formation of waves in traffic flow, and has helped to popularise it. A more recent experiment under similar conditions has shown that a single vehicle with particularly stable dynamics allows the waves to dissipate. This field of research remains particularly active (see section 4 below).
Figure 3. Trajectories in the experiment of Sugiyama et al. with 22 vehicles on a circuit of length 230 meters starting from an uniform configuration [114]. After a while, a stop-and-go wave appears, which causes a decrease in the average speed and an increase in the speed standard deviation.
@@ 3.1.2. Collective stability analysis
The formation of stop-and-go waves and phase separation from laminar traffic to oscillatory dynamics was theorised in the 2010s, mostly using linear stability analysis [127, 96, 139]. Local stability analysis considers a platoon of vehicles following a leader. Stability conditions are usually derived using Laplace transform and transfer function. Global (or string or collective) stability analysis considers a flow of vehicles with periodic boundary conditions. In general, global stability conditions are more restrictive than local ones, since global analysis takes into account advective or convective perturbations that vanish locally. The stability conditions for infinite systems are even more restrictive. The stability of finite periodic systems can be derived using matrix spectral analysis, while for infinite systems an exponential ansatz is generally used. The latter is equivalent to performing a Laplace transformation in time and a discrete Fourier transformation in space [96].
@@
see, e.g., [123], [127, Chap. 15] and general stability conditions for complex polynomials in [39, Th. 3.2] which are a generalisation of the Hurwitz conditions. The conditions are sufficient for any finite system and become exact for infinite systems. For example, the linear stability condition is
@@
for the FVD model (8), while the ATG model (10) is systematically linearly stable provided that $\tau,T>0$, hence its prevalence for automated driving systems [61].
@@ 3.2.1. Empirical observations in pedestrian and other particle systems
Figure 4. Trajectories of pedestrians in a counterflow experiment [14]. The blue trajectories are pedestrians moving from right to left, while the grey trajectories are pedestrians moving from left to right. We can see the formation of lanes moving in opposite directions.
Counterflows, where two groups of pedestrians move in opposite directions, often lead to the formation of lanes of pedestrians walking in the same direction. Lane formation in pedestrian dynamics has been observed since the 1960s [95, 89] and is studied in experiments (see Fig. 4 and [14]). It may also be related to the stripes formed at intersections between pedestrian streams [4]. In fact, the phenomenon of lane formation is widely encountered in discrete systems in which entities move in two opposite directions [33]. For example, lane formation occurs in passive matter, e.g., driven mixtures of oppositely charged colloids [135] or driven plasmas [115, 105], and in active matter, such as in microswimmers [64], bacterial populations [109], or social insects such as ants [26]. Nowadays, laning phenomenon is recognised as a fundamental collective property of particle and self-avoiding agent systems driven by a mixture of two potentials [28, 101].
@@ 3.2.2. Modelling lane formation
Lane formation in pedestrian and other mixed-particle models is difficult to analyse mathematically and is usually studied by simulation or using continuous appraoches (see, e.g., [4]). In fact, most of two-dimensional pedestrian models are able to reproduce laning in counterflow. Pioneering simulation results were obtained in the 2000s using the social force model [48], basic cellular automata [12], or, a few years later, velocity obstacle models [45]. The formation of lanes can also be explained by socio-psychological factors of group membership and social identity [117, 138]. In [60], lane formation is even observed for mixed flows moving in the same direction but at different speeds. However, some experiments with pedestrians walking at different speeds did not provide empirical evidence of lane formation [40].
In fact, the occurrence of lane formation is observed also in systems of inanimate particles. For example in [115], the volume exclusion and collisions due to the motion in different directions in driven binary complex plasma, combined with joint pushing of particles moving in the same direction, induce a segregation into lanes. Thus, any minimal models for pedestrian dynamics representing volume exclusion, such as force models with repulsive interactions, are able to reproduce lane formation. This example indicates that lane formation itself does not need social or psychological explanations, even if the steering process of humans in such a situation is based on the impressive cognitive and biomechanical abilities of the human body. These aspects are discussed in more detail in [111].
In general, lane formation in pedestrian counterflow depends on density. At low densities, it may be advantageous to avoid rather than follow, thus limiting lane formation. Conversely, at high densities, exclusion constraints may prevent the formation of lanes. However, there is no empirical evidence for a transition, although experiments have shown that counterflow can give rise to gridlock at densities greater than 3.5 persons per square metre [146]. In the models, these density thresholds above which lanes no longer emerge depend on the form of the model in addition to the initial configuration [142, 121]. They remain an open rearch question. Another factor that leads to counterflow gridlock is the introduction of noise into pedestrian motion models. This phenomenon is known in the literature as freezing by heating [49]. Even in simple deterministic models, a transition to lane formation occurs as the driven force increases relative to the interacting repulsive forces [124].
@@ 3.3. Far-sighted anticipation in local navigation
Rather than lanes arising from oppositely-directed flows, let us now inspect the response of a static crowd to a moving ‘obstacle’. For this purpose, we focus on an experiment presented in [93] that highlights the long-term anticipatory abilities of pedestrians to avoid an intruder, and how mean-field games can be used to model this type of behaviour.
@@ 3.3.1. Empirical evidence for long-term anticipation
That pedestrians or cars plan their motion tens of minutes or even hours ahead should not come as a surprise. Unlike physical particles, their local navigation is steered towards a target destination set at larger modelling scales. The existence of a desired direction, imparted to agents at the higher tactical level, makes space anisotropic for them (i.e., not all directions are equivalent), irrespective of their interactions with the rest of the crowd, in contrast to boids [103] or simple swarming models [134] for which directional preferences only result from interactions.
What may be less obvious is that even the local, operational interactions with their neighbours and the built environment may involve some degree of planning. Direct visualisation of the simulation output of simple force-based models and the frequent near-collisions they imply evince that distance-based interaction forces are not sufficient to render realistic collision avoidance. Taking into account the relative velocities of the agents is thus necessary. Evaluating distances via this relative velocity, i.e., in terms of the imminence of a collision risk, gives rise to the well-known notion of anticipated times to collision (TTC). Experimental observations (collected in real life or in virtual reality) as well as empirical data show the central role played by the TTC for collision avoidance and the spatial arrangement of crowds [58, 100, 82, 25]. Interactions between people may start at a TTC of several seconds, when they are more than half a dozen meters away from each other.
Interestingly, anticipation may even need to be pushed beyond the nearest collision to describe operational pedestrian dynamics [136]. In particular, a crowd’s response to an intruder’s crossing [93] (by transverse moves opening a depleted tunnel ahead of the intruder, see Fig. 5) appears to elude state-of-the-practice models based (or not) on TTC [11, 13]. The observed response could actually be understood by assuming that pedestrians anticipate much beyond the next collision, and actually beyond the passage of the moving obstacle, as we shall see.
Figure 5. Experiment [93] in which a dense crowd is crossed by an cylindrical intruder (Left). Pedestrians anticipate the passage of the cylinder as seen from the void in front of the cylinder, and temporarily accept the discomfort of the dense regions on both sides of the cylinder (Right, from [13]), because they know it will not last.

### v2

@@ 2.2.3. Hybrid models
Many pedestrian models are hybrid approaches combining different concepts and mechanisms from first and second order models such as social force, optimal velocity (OV), or optimisation and collision avoidance techniques. For example, some extensions of force-based models are inspired by swarming approaches [helbing2000simulating, lakoba2005modifications], where the desired velocity $\vec{v}_{0}$ in (11) is mitigated with the average velocity of the neighbouring pedestrians using
@@
to address mass collective behaviour. In the same vein, the pedestrian velocity is relaxed to a combination of the desired velocity and a superposition of distance-based OV models [nakayama2005instability, nakayama2008effect], generalising concepts of car-following models [bando1995dynamical, bando1998analysis, jiang2001full] to two dimensions
@@
Another hybrid two-dimensional OV pedestrian model relies on first-order dynamics [tordeux2016collision]. As in the modelling framework introduced by Maury and Venel in [maury2008mathematical, maury2011discrete], this model constrains pedestrian velocities to remain in admissible collision-free configurations [tordeux2016collision]. However, the speed is regulated according to the distance to the next pedestrian in the direction of motion using an OV function, while the direction results from a superposition of distance-based repulsion with the neighbours as in force-based models. An extension of this collision-free velocity-based model relaxes the direction, making it a second-order model [xu2019generalized]. In the Gradient Navigation model introduced in [dietrich2014gradient], the speed is relaxed to the OV function, while the direction results from a the desired velocity and a superposition of distance-based repulsion with the neighbours as in force-based models
@@
where $C_{1}$ and $C_{2}$ are two normalisation constants while $V(\rho_{n})$ is the OV depending on the local density $\rho_{n}$. A similar approach is given in [moussaid2011simple] with an heuristic model without social interaction and where, as in the collision-free velocity-based model [tordeux2016collision], the OV depends on the distance to the nearest pedestrian in front.
The concept of anticipatory collision avoidance in force-based models introduced in [zanlungo2011social] is deepened in a recent hybrid force-based model [karamouzas2014universal], where the interaction force is derived directly from the potential
@@
The TTC $\tau_{n,m}$ is the time before a collision occurs between the $n$-th and $m$-th pedestrians (i.e., $|\Delta\dot{\vec{x}}_{n,m}|<\ell$), assuming that their velocities are constant. In [lu2020pedestrian], the force is based on similar collision anticipation mechanisms. A general hybrid modelling framework first introduced in the early 2000s [hoogendoorn2003simulation] assumes pedestrian dynamics based on a decisional layer, resulting from an optimal control problem, and a mechanical layer modelling physical exclusion force and other inertia phenomena
(25) $\tilde{\vec{v}}_{0}^{n}(t)=\text{arg}\hskip-2.84526pt\min_{\vec{v}_{n}\in\mathbb{R}^{2}}\mathcal{E}_{\vec{x}(t)}^{\vec{v}_{0}}(\vec{v}_{n}),\qquad\ddot{\vec{x}}_{n}(t)=\frac{1}{\tau}\big[\tilde{\vec{v}}_{0}^{n}(t)-\dot{\vec{x}}_{n}(t)\big]+F^{\text{M}}_{n}(t),$
where $\mathcal{E}$ is an energy function involving several concepts, while $F^{\text{M}}_{n}(t)$ are the mechanical force on the $n$-th pedestrian at time $t$. In the ANticipatory Dynamic Algorithms (ANDA) [echeverria2023body], the energy function consists of four main components:
@@
    Anticipation energy based on the TTC potential (23)-(24) introduced in [karamouzas2014universal].
@@ 3. Modelling collective dynamics
Traffic and pedestrian flows describe many coordinated collective dynamics [helbing2001traffic, chowdhury2000statistical, castellano2009statistical, boltes2018empirical] A typical example in traffic flow is stop-and-go dynamics on highways where jam waves resulting from instability phenomena [orosz2010traffic, wilson2011car], also called jamitons [flynn2009self], propagate upstream in the flow without any apparent external cause. Typical examples for pedestrian dynamics are lane formation for counterflow, jamming and clogging at bottlenecks, oscillation and flow intermittency for counterflow at bottlenecks, or stripe formation for crossing flow [helbing2005self, boltes2018empirical]. Most of the aforementioned models describe these phenomena, with realistic features for fine-tuning the parameters. However, conventional models fall short of replicating more complex situations involving long-term anticipation (on the order of several seconds) or socio-psychological aspects. A typical example is the formation of a clear lane in pedestrian crowds to facilitate the passage of a vehicle, typically an ambulance [MadrasAmbulance], which relies on early anticipation. Another example in the load balancing of pedestrians over several exits in case of normal evacuation. Here, the distance between the exits requires pedestrians to anticipate for long periods of time. These different collective phenomena and related modelling approaches are discussed in more detail in the following sections.
@@ 3.1.1. Empirics of stop-and-go waves in traffic flow
Stop-and-go dynamics are currently observed in road traffic flow. They are also found in single-file pedestrian movements and in bicycle flow [zhang2014universal]. In fact, stop-and-go and accordion traffic now commonly denotes congested traffic. In addition to its scientific interest, the jerky dynamics of congested traffic is a major safety concern. It also has a negative impact on the environment. Indeed, the acceleration and deceleration phases of stop-and-go traffic dynamics cause an increase in fuel consumption and pollutant emissions [li2014stop].
The terminology of stop-and-go waves can be traced back to the late 1960s [duckstein1967control]. A related word is phantom jam [sugiyama2008traffic], as the waves seem to appear for no apparent reason. This was demonstrated in the late 2000s with an experiment involving 22 vehicles on a 230-metre circuit [sugiyama2008traffic] (see Fig. 3). Starting from a homogeneous configuration with equal distance gaps, a stop-and-go wave emerged after a while, with jerky vehicular dynamics not caused by any infrastructure peculiarity. This experiment shed new light on the formation of waves in traffic flow, and has helped to popularise it. A more recent experiment under similar conditions has shown that a single vehicle with particularly stable dynamics allows the waves to dissipate. This field of research remains particularly active (see section 4 below).
Figure 3. Trajectories in the experiment of Sugiyama et al. with 22 vehicles on a circuit of length 230 meters starting from an uniform configuration [sugiyama2008traffic]. After a while, a stop-and-go wave appears, which causes a decrease in the average speed and an increase in the speed standard deviation.
@@ 3.1.2. Collective stability analysis
The formation of stop-and-go waves and phase separation from laminar traffic to oscillatory dynamics was theorised in the 2010s, mostly using linear stability analysis [treiber2013traffic, orosz2010traffic, wilson2011car]. Local stability analysis considers a platoon of vehicles following a leader. Stability conditions are usually derived using Laplace transform and transfer function. Global (or string or collective) stability analysis considers a flow of vehicles with periodic boundary conditions. In general, global stability conditions are more restrictive than local ones, since global analysis takes into account advective or convective perturbations that vanish locally. The stability conditions for infinite systems are even more restrictive. The stability of finite periodic systems can be derived using matrix spectral analysis, while for infinite systems an exponential ansatz is generally used. The latter is equivalent to performing a Laplace transformation in time and a discrete Fourier transformation in space [orosz2010traffic].
@@
see, e.g., [tordeux2012linear], [treiber2013traffic, Chap. 15] and general stability conditions for complex polynomials in [Frank1946OnTZ, Th. 3.2] which are a generalisation of the Hurwitz conditions. The conditions are sufficient for any finite system and become exact for infinite systems. For example, the linear stability condition is
@@
for the FVD model (8), while the ATG model (10) is systematically linearly stable provided that $\tau,T>0$, hence its prevalence for automated driving systems [khound2023extending].
@@ 3.2.1. Empirical observations in pedestrian and other particle systems
Figure 4. Trajectories of pedestrians in a counterflow experiment [cao2017fundamental]. The blue trajectories are pedestrians moving from right to left, while the grey trajectories are pedestrians moving from left to right. We can see the formation of lanes moving in opposite directions.
Counterflows, where two groups of pedestrians move in opposite directions, often lead to the formation of lanes of pedestrians walking in the same direction. Lane formation in pedestrian dynamics has been observed since the 1960s [oeding1963verkehrsbelastung, navin1969pedestrian] and is studied in experiments (see Fig. 4 and [cao2017fundamental]). It may also be related to the stripes formed at intersections between pedestrian streams [bacik2023lane]. In fact, the phenomenon of lane formation is widely encountered in discrete systems in which entities move in two opposite directions [dzubiella2002lane]. For example, lane formation occurs in passive matter, e.g., driven mixtures of oppositely charged colloids [vissers2011lane] or driven plasmas [sutterlin2009dynamics, sarma2020lane], and in active matter, such as in microswimmers [kogler2015lane], bacterial populations [shimaya2019lane], or social insects such as ants [couzin2003self]. Nowadays, laning phenomenon is recognised as a fundamental collective property of particle and self-avoiding agent systems driven by a mixture of two potentials [cristin2019general, reichhardt2018laning].
@@ 3.2.2. Modelling lane formation
Lane formation in pedestrian and other mixed-particle models is difficult to analyse mathematically and is usually studied by simulation or using continuous appraoches (see, e.g., [bacik2023lane]). In fact, most of two-dimensional pedestrian models are able to reproduce laning in counterflow. Pioneering simulation results were obtained in the 2000s using the social force model [helbing2000simulating], basic cellular automata [burstedde2001simulation], or, a few years later, velocity obstacle models [guy2010pledestrians]. The formation of lanes can also be explained by socio-psychological factors of group membership and social identity [templeton2020placing, wijermans2022towards]. In [khelfa2022heterogeneity], lane formation is even observed for mixed flows moving in the same direction but at different speeds. However, some experiments with pedestrians walking at different speeds did not provide empirical evidence of lane formation [fujita2019traffic].
In fact, the occurrence of lane formation is observed also in systems of inanimate particles. For example in [sutterlin2009dynamics], the volume exclusion and collisions due to the motion in different directions in driven binary complex plasma, combined with joint pushing of particles moving in the same direction, induce a segregation into lanes. Thus, any minimal models for pedestrian dynamics representing volume exclusion, such as force models with repulsive interactions, are able to reproduce lane formation. This example indicates that lane formation itself does not need social or psychological explanations, even if the steering process of humans in such a situation is based on the impressive cognitive and biomechanical abilities of the human body. These aspects are discussed in more detail in [sieben2017collective].
In general, lane formation in pedestrian counterflow depends on density. At low densities, it may be advantageous to avoid rather than follow, thus limiting lane formation. Conversely, at high densities, exclusion constraints may prevent the formation of lanes. However, there is no empirical evidence for a transition, although experiments have shown that counterflow can give rise to gridlock at densities greater than 3.5 persons per square metre [zhang2012ordering]. In the models, these density thresholds above which lanes no longer emerge depend on the form of the model in addition to the initial configuration [xu2021anticipation, tordeux2018mesoscopic]. They remain an open rearch question. Another factor that leads to counterflow gridlock is the introduction of noise into pedestrian motion models. This phenomenon is known in the literature as freezing by heating [helbing2000freezing]. Even in simple deterministic models, a transition to lane formation occurs as the driven force increases relative to the interacting repulsive forces [tordeux2023multi].
@@ 3.3. Far-sighted anticipation in local navigation
Rather than lanes arising from oppositely-directed flows, let us now inspect the response of a static crowd to a moving ‘obstacle’. For this purpose, we focus on an experiment presented in [nicolas2019mechanical] that highlights the long-term anticipatory abilities of pedestrians to avoid an intruder, and how mean-field games can be used to model this type of behaviour.
@@ 3.3.1. Empirical evidence for long-term anticipation
That pedestrians or cars plan their motion tens of minutes or even hours ahead should not come as a surprise. Unlike physical particles, their local navigation is steered towards a target destination set at larger modelling scales. The existence of a desired direction, imparted to agents at the higher tactical level, makes space anisotropic for them (i.e., not all directions are equivalent), irrespective of their interactions with the rest of the crowd, in contrast to boids [reynolds1987flocks] or simple swarming models [vicsek1995novel] for which directional preferences only result from interactions.
What may be less obvious is that even the local, operational interactions with their neighbours and the built environment may involve some degree of planning. Direct visualisation of the simulation output of simple force-based models and the frequent near-collisions they imply evince that distance-based interaction forces are not sufficient to render realistic collision avoidance. Taking into account the relative velocities of the agents is thus necessary. Evaluating distances via this relative velocity, i.e., in terms of the imminence of a collision risk, gives rise to the well-known notion of anticipated times to collision (TTC). Experimental observations (collected in real life or in virtual reality) as well as empirical data show the central role played by the TTC for collision avoidance and the spatial arrangement of crowds [karamouzas2014universal, pfaff2018avoidance, meerhoff2018guided, cordes2024dimensionless]. Interactions between people may start at a TTC of several seconds, when they are more than half a dozen meters away from each other.
Interestingly, anticipation may even need to be pushed beyond the nearest collision to describe operational pedestrian dynamics [wagoum2017understanding]. In particular, a crowd’s response to an intruder’s crossing [nicolas2019mechanical] (by transverse moves opening a depleted tunnel ahead of the intruder, see Fig. 5) appears to elude state-of-the-practice models based (or not) on TTC [bonnemain2023pedestrians, butano_a_u2024a]. The observed response could actually be understood by assuming that pedestrians anticipate much beyond the next collision, and actually beyond the passage of the moving obstacle, as we shall see.
Figure 5. Experiment [nicolas2019mechanical] in which a dense crowd is crossed by an cylindrical intruder (Left). Pedestrians anticipate the passage of the cylinder as seen from the void in front of the cylinder, and temporarily accept the discomfort of the dense regions on both sides of the cylinder (Right, from [butano_a_u2024a]), because they know it will not last.